OpenStudy (anonymous):

I'm supposed to find the radius of convergence and radius of convergence. \[\sum_{n=0}^{\infty} \frac{(-1)^{n}x^{n}}{n+1}\]

OpenStudy (anonymous):

I started with the ratio test

OpenStudy (anonymous):

this is what I came up with \[\lim_{n \rightarrow \infty} \left|x\right| \frac{n+1}{n+2}\]

OpenStudy (anonymous):

=x

OpenStudy (anonymous):

right? \[\lim_{n \rightarrow \infty} \left|x\right| \frac{n/n+1/n}{n/n+2/n}\]

OpenStudy (anonymous):

Seems right.

OpenStudy (anonymous):

where from here?

OpenStudy (anonymous):

The ratio test states that, if we call the limit L, so L=|x| here, if L<1 the series converges if L>1 the series diverges if L=1 inconclusive

OpenStudy (anonymous):

so as long as x<1

OpenStudy (anonymous):

To be precise: |x|<1, so -1<x<1

OpenStudy (anonymous):

just curious why -1<x ?

OpenStudy (anonymous):

|x|<1 is equivalent to -1<x<1 if x=-2 |x|=|-2|=2 is not smaller than 1.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

You can see that the series diverges for a large negative x, -10 for example: 5+100/3+1000/4+...

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Also the series might converge for x=1 or -1 as well, the test is inconclusive there.

OpenStudy (anonymous):

something about harmonic series and alternating harmonic series?

OpenStudy (anonymous):

yes